OpenStudy (anonymous):

11 mins is what fractional part of 11 hours?

OpenStudy (anonymous):

Convert 11 hours to the same units of minutes-->11 hours * 60 mins/hr=660 mins. Then, take 11/660 to get the fraction that 11 is of 11 hours = 1/60

myininaya (myininaya):

11/60 11/(11*60)=1/60

myininaya (myininaya):

i multiply by 1/11 because we have 11 hours not just the 1